OpenStudy (anonymous):

A ball is thrown straight up in the air at a velocity of 24 feet/second. The height of the ball at t Seconds is represented by the equation h=24t-16t^2. find the maximum height and the time at which it will be there? can someone walk me through this?

OpenStudy (anonymous):

Flvs?

OpenStudy (anonymous):

any ways the way you solve this is by they equation f(t) = -16t^2 + vt + s

OpenStudy (anonymous):

This should be on the question

OpenStudy (anonymous):

It says use the equation that I gave to solve?

OpenStudy (anonymous):

I dont see the equation

OpenStudy (anonymous):

Please post it so i can help

OpenStudy (anonymous):

look at the pic

OpenStudy (anonymous):

Ok got it thank you

OpenStudy (anonymous):

Go ahead and graph that in something like GeoGebra or Wolfram Alpha

OpenStudy (anonymous):

k

OpenStudy (anonymous):

You will get the anwers in bot of the because it tells you where the max and min is with a picture of the graph which is really helpfull

hero (hero):

Find h'

hero (hero):

then set h' = 0. Then solve for t

OpenStudy (phi):

**what are the white people when you look at viewing ? People not logged in

OpenStudy (anonymous):

ok sorry my internet is being slow

OpenStudy (phi):

If this is algebra (not calculus), you can find the peak of the parabola. It is the vertex of the parabola

OpenStudy (anonymous):

so with this i just do the middle term. and yeah its algebra but I couln't figure out the labeling for the vertex equation

OpenStudy (phi):

write the equation in standard form (x^2 term is first)

OpenStudy (phi):

\[ h= -16t^2+ 24t + 0 \] match to y = a x^2 + b x + c

OpenStudy (anonymous):

ok got that

OpenStudy (phi):

vertex is at t= -b/(2a)

OpenStudy (anonymous):

-.75?

OpenStudy (phi):

a= -16 , b= 24 -b/(2a) is -24/-32 = + 24/32 = ¾ = 0.75 so you were close.

OpenStudy (anonymous):

ok

OpenStudy (phi):

you should notice from the graph that the vertex does not occur at a negative t. that is a clue to double check what you did.

OpenStudy (anonymous):

so i thats my line of symmetry?

OpenStudy (phi):

yes. notice that 0.75 is exactly ½ between the "zeros" at x=0 and x=1.5 See graph

OpenStudy (anonymous):

great thanks!